

********************************************************************
UARS WEEKLY STATUS REPORT - 17 SEP 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues normal operations. Complete daily solar data sets
have been collected this week along with continued stellar calibrations.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control. While
under PMON control, the PEM instrument computed energy deposition from its
particle spectrometers.
The FOT and the PEM team are updating the PEM activation sequence which
effects the PEM/HEPS (> 30 keV particles). Current activation sequence
reduces the activation time by about 1 minute. Additional analization shows
that the activation sequence may be reduced by about another minute. PEM
PMON activation is set for the beginning of power orbit day. The beginning of
power orbit day is defined as when:
array current greater than 5 amps, and
array voltage greater than 53.5 volts, and
any battery current greater than 2.0 amps
The end of power orbit day (PEM deactivation) is defined as when:
time in power orbit day greater than 4 minutes, and
array current less than 9 amps, and
all battery currents less than 0.2 amps
Power orbit day periods vary depending upon beta angle:
about 39-40 minutes at yaw
about 50-51 minutes at max beta
Current activation takes about 7.5 min for full activation. During the
activation sequence, various components of PEM become active at all different
times, beginning with the PEM magnetometer (at about on activation time) and
ending with the HEPS-2 components (7.5 min later).
***********
HALOE
***********
HALOE continues to have successful operations. HALOE is scheduled to turn
off at the end of day September 21, 1999.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
18 days CLA (7.0) PROD (L1-L3) (473R-256C-15F=202L)
0 days GPI (4.05) PROD (CLAES L3B) (191R-164C-2F=25L)
0 days WIN (5.11A)TEST (L1-L3) (3R-0C-0F=3L)
The CDHF will load July 1999 current day DAAC data in September 1999.
CDHF personnel continued to draft procedures for the new Alpha 4000 and
HSM systems.
CDHF pending software upgrades: none
***********
FOT
***********
REPORTING PERIOD
This report covers 4 September, 1999 (Orbit 43612, GMT Day 247) through 10
September 1999 (Orbit 43716, GMT Day 253).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 247/0404 43614 Off
PEM 252/2316 43701 Off
WINDII 253/0100 43702 Off
The beta angle ranged between the angles of 49.5 degrees to 47.3 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
-2 degrees on 24 September 1999 (DOY 267).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.7 to 1.0 degrees C. Additional battery performance data for each
battery on 03 September 1999 and 10 September 1999 are provided in Appendix D.
The clock error ranged from -10.3 to 12.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 03
% PEM = 02
% WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 05
Zero (0) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/242 - 2287.496953
% Transponder B: 99/243 - 2287.499030
GROUND SYSTEM OPERATION
Three (5) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/252 show the observatory data loss to be 24 hours, 55
minutes, 09.428 seconds (an increase of 18.912 seconds since the last report
period). This is a 0.0357 percent data loss which equals a 99.9643 percent
data capture for the mission.
No increase to report for this week.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
PEM requested that to have the timing in their turn on sequence in yet another
area tightened up. This is expected to be accomplished within the next few
weeks.
YEAR 2000 OPERABILITY SUPPORT
End-to-end Y2K testing continued this past week. Much of the effort this past
week was spent in debugging and configuring the scenario into the simulator.
Dennis Dixon was called for consultation and support. Testing was postponed
until next week while a Mean Element report from FDF was requested.
The FOT is waiting for a valid 8-week PSAT file from FDF to be used for Y2K
testing. The previous PSAT file that was delivered was actually only a 4-day
file, not an 8-week one. Without a complete set of 8-week files, the UARS CMS
is unable to perform all of the necessary mission planning and load building
processes required for Y2K end-to-end testing. In the meantime, the FOT has
generated a "test" Y2K ATC load by editing an existing 1999 load, and
performing a universal substitution edit command on the year, changing it from
1999 to 2000. This ATC load has not been tested on the simulator as of yet.
MEETINGS
The monthly meeting was held this week. There were no new topics mentioned at
the meeting, but we did receive a deadline of next month for the Leonid Plan.
Three members of the UARS/TOMS staff attended a security briefing given by Andy
Kochis, the CSOC Security Manager. We were briefed on all security issues that
Goddard is trying to crack down on such as access to government equipment and
password criteria. This meeting lasted from 8:30 am to 12:00 pm.
FUTURE SIGNIFICANT EVENTS
A meeting is planned for in September to discuss the UARS Safehold
configuration and concerns for the MPS and its batteries. This meeting will be
a closed meeting with only the FOT and specialized engineering support
attending.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 10 September 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
43614 247/0404 99-314 PEM Axis Power ON
43634 248/1136 99-315 OBC Clk Adj fr 20.6 to 0.00 msec/day
43662 250/0826 99-316 OBC Clk Adj fr 0.00 to 20.6 msec/day
43696 252/1532 99-317 OBC Clk Adj fr 20.6 to 0.00 msec/day
43701 252/2316 99-318 PEM AXIS Power OFF
43702 253/0100 99-319 WINDII Power ON
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 10 September 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI Quiet Bus Current Monitor 43705-6
NBRECBI Y-HI NBTR B Current 43705-6
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 10 September 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---------------------- -------------------------
43634 248/11:41 (05 Sept 99) from 20.6 to 00.0 msec/day
43662 250/08:29 (07 Sept 99) from 00.0 to 20.6 msec/day
43696 252/15:32 (09 Sept 99) from 20.6 to 00.0 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 10 September 1999
03 Sept 1999 - GMT Day 232 Beta = 49.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- -------- --------- ----- ------- ------
1 -112.0/-134.4 -3.05/-4.31 +0.0/+0.0 16.5 0.000 0.0
2 +39.2/+0.0 +5.47/+4.51 +26.4/-15.2 25.3 1.019 21.4
3 +33.6/-11.2 +4.51/+3.87 +25.6/-15.2 25.1 1.023 21.1
10 Sept 1999 - GMT Day 253 Beta = 47.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- -------- --------- ---- ------- ------
1 -112.0/-134.4 -3.05/-4.31 +0.0/+0.0 16.5 0.000 0.0
2 +39.2/-11.2 +5.79/+4.19 +27.2/-16.0 24.5 1.024 21.0
3 +44.8/-33.6 +4.83/+3.23 +26.8/-16.0 24.3 1.024 20.8
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 10 September 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-051 FOT 43617 247/0817 231310 No GCMR capability
99-051 FOT 43644 249/0409 N/A TAC-6 Crashed
99-052 FOT 43655 249/2119 N/A TAC-4 Crashed
99-053 FOT 43664 250/1248 N/A TAC-8 Crashed
99-054 FOT 43668 250/1848 N/A TAC-8 Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-033 FOT 42856 196/1323 N/A TMON #1 trip
99-040 FOT 43187 218/1634 21201 NASCOM problem at WSGT
99-043 FOT 43458 236/1855 21265 Late Acq / sched.conflict
99-044 FOT 43543 242/1114 N/A AP Crash - NIU down
99-045 FOT 43574 244/1236 N/A AP Crashed-switched Pack
99-046 FOT 43576 244/1607 N/A TAC 4 Crashed
99-050 FOT 43617 247/0817 231310 No GCMR capability
99-051 FOT 43644 249/0409 N/A TAC-6 Crashed
99-052 FOT 43655 249/2119 N/A TAC-4 Crashed
99-053 FOT 43664 250/1248 N/A TAC-8 Crashed
99-054 FOT 43668 250/1848 N/A TAC-8 Crashed
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None.
SEVNTFW99/37